4) For a Given Operational Amplifier Please Determine the following Quantities, Amid Voltage Gain 450 f1 125 Hz f2 625 Hz Voltage Gain 60 Hz a) 188.66 b) 28.80 c) 194.72 d) 13.77 Voltage Gain 1250 Hz a) 201.24 b) 70.70 c) 154.77 d) 200.60 Voltage Gain @ f1 and f2 a) 311.23 b) 344.98 c) 318.15 d) 240.66
Answer #1
  1. At 60 Hz i.e. for frequency below the lower cut-off frequency, the gain is-
    450 = 194.8 160 Thus option C is the correct answer.

  2. At 1250 Hz i.e. for frequency higher tha the higher cut-off frequency, the gain is-
    G=\frac{450}{\sqrt{1+(\frac{1250}{625})^{2}}} = 201.25   Thus option A is the correct answer.

  3. At cut-off frequency the gain is 0.707 times of maximum value. So G = 0.707*450 = 318.15. Thus option C is the correct answer.